Calculate the rate at which a blackbody at temperature 300°C emits radiant energy. The blackbody is a sphere of diameter 20cm.
1
Expert's answer
2020-08-17T08:57:54-0400

According to the Stefan–Boltzmann law, the total energy radiated per unit surface area of a black body across all wavelengths per unit time is:


"j = \sigma T^4"

where "\sigma = 5.67\times 10^{-8} Wm^{-2}K^{-4}" is the Stefan–Boltzmann constant and "T = 300\degree C = 573K" is the a blackbody temperature.

In order to obtain the radiation rate one should multiply "j" by the area of blackbody surface:


"P = A\cdot j = \pi d^2j"

where "d = 20cm = 0.2m" is the diameter of the sphere. Thus, obtain:


"P = \pi d^2j = \pi d^2\sigma T^4 = \pi\cdot 0.2^2\cdot 5.67\times 10^{-8}\cdot 573^4 \approx 768.1\space W"

Answer. 768.1 W.